# From Video to EEG: Adapting Joint Embedding Predictive Architecture to Uncover Visual Concepts in Brain Signal Analysis

### v2

@@ Abstract
EEG signals capture brain activity with high temporal and low spatial resolution, supporting applications such as neurological diagnosis, cognitive monitoring, and brain-computer interfaces. However, effective analysis is hindered by limited labeled data, high dimensionality, and the absence of scalable models that fully capture spatiotemporal dependencies. Existing self-supervised learning (SSL) methods often focus on either spatial or temporal features, leading to suboptimal representations. To this end, we propose EEG-VJEPA, a novel adaptation of the Video Joint Embedding Predictive Architecture (V-JEPA) for EEG classification. By treating EEG as video-like sequences, EEG-VJEPA learns semantically meaningful spatiotemporal representations using joint embeddings and adaptive masking. To our knowledge, this is the first work that exploits V-JEPA for EEG classification and explores the visual concepts learned by the model. Evaluations on the publicly available Temple University Hospital (TUH) Abnormal EEG dataset show that EEG-VJEPA outperforms existing state-of-the-art models in classification accuracy.Beyond classification accuracy, EEG-VJEPA captures physiologically relevant spatial and temporal signal patterns, offering interpretable embeddings that may support human-AI collaboration in diagnostic workflows. These findings position EEG-VJEPA as a promising framework for scalable, trustworthy EEG analysis in real-world clinical settings
@@ 1 Introduction
    We present EEG-VJEPA, a novel self-supervised model based on the joint embedding predictive architecture that incorporates the spatial and temporal attributes of EEG signals. We treat EEG signals as video-like sequences, enabling the application of V-JEPA’s saptiotemporal masking strategy.
@@
    We show that self-supervised learning with the VJEPA strategy provides a major improvement compared to EEG2REP [19], LaBraM [20], and the model trained with the contrastive learning approach [21] on the clinically relevant TUAB dataset. We attribute this improvement to the high semantic saptiotemporal representations learned in the latent space.
@@
The remainder of the paper is structured as follows: Section 2 provides an overview of the SSL literature. In Section 3, we present the methodology of our work. Section 4 describes the experimental settings, highlighting the implementation details and results. In Section 5, we discuss our findings and implications. Finally, in Section 6, we conclude the paper.
@@ 2 Related Work
A significant amount of progress has been made in the processing and creation of useful representations of EEG data, focused on feature extraction, pre-training, predictive modeling, and downstream task performance. Early works transformed time-series EEG data into frequency domains and combined statistical and spectral features for classification using models such as SVMs [23]. Other studies leveraged joint representations from time and frequency domains via convolutional layers [24] or multi-view approaches such as wavelet transforms and attention-based gated recurrent units [25]. Incorporating brain connectivity measures like phase lag index and power spectral density has also enhanced predictive modeling [26].
Self-supervised learning (SSL) methods have shown promise in addressing the scarcity of labeled EEG data. Multi-view SSL approaches have fused diverse representations, such as time, frequency, and spatial features, for tasks like motor imagery classification and seizure prediction [27, 24]. A work by [21] developed a self-supervised approach for EEG representation learning using multi-view contrastive learning. The method utilized both time and frequency domain views of EEG data with data augmentation. [28] introduced EEGformer, a model combining convolutional and Transformer layers to capture temporal and regional EEG features. These methods demonstrated strong performance in emotion recognition and seizure detection. Recent works also highlight the potential of large-scale pre-training for EEG. Chen et al. [29] developed a 1.7TB EEG foundation model using a vector-quantized Transformer, offering interpretable representations with strong transfer learning capabilities. The Joint Embedding Predictive Architecture (JEPA), a self-supervised framework for learning semantic representations by predicting missing information in latent space without data augmentation, was introduced by  [30]. This approach has been adapted for various data types, including EEG signals. For instance, Guetschel et al. proposed Signal-JEPA, which incorporates a domain-specific spatial block masking strategy to enhance cross-dataset transfer in EEG signal processing [31]. However, these approaches often fail to capture the long-range temporal context critical for EEG analysis. Inspired by V-JEPA [22], we extend this architecture to EEG by treating signals as video-like sequences.
@@ 3 Methodology
We process the EEG sequence by first dividing each frame into non-overlapping spatiotemporal volumes or patches. Each patch captures a small region of the EEG signal across several channels and time points. We then apply a patch embedding layer, using a combination of convolutions and linear projections, to map each patch to a fixed-dimensional embedding vector.
The x-encoder processes the sequence of patch embeddings after masking a portion of the patches. The masking strategy is crucial for creating the prediction task. We adapt V-JEPA’s multi-block masking approach, which involves masking large, spatially contiguous blocks that span the entire temporal dimension of the sequence. This strategy encourages the model to learn long-range spatiotemporal dependencies. During masking the aspect ratio for the sampled blocks is randomly chosen in a certain range (see Table 1).
@@ 4.1 Experimental Setting
Datasets: We use two publicly available datasets, namely TUAB [18] and NMT [32], to train and evaluate our model.
@@
[…]
@@
Following the literature, we use the following naming scheme for the ViT models: ViT-Base (ViT-B, $L=12$, $NH=12$, $d=768$) with 85 M parameters, ViT-Medium (ViT-M, $L=12$, $NH=6$, $d=384$) with 21 M parameters and ViT-Small (ViT-S, $L=12$, $NH=3$, $d=192$) with 5 M parameters, where $L$ is the number of transformer layers, each with a self-attention block of $NH$ heads. To indicate the tubelet size while patching, we extend the naming scheme, for example, ViT-B/$4\times 30\times 2$ denotes a ViT-Base backbone with the tubelet size of ($h\times w\times t=4\times 30\times 2$). we tested various ViT encoder sizes (base, medium, small) to evaluate the impact of model capacity on representation learning; refer to Table 1 for more details. In all our experiments, the tubelet height ($h$) and width($w$) differ due to the shape of our EEG input signal height and width ($19\times 500$). We also probe with the time dimension $t$ while patching.
Pre-training: We combine two publicly available datasets, namely TUH [18] and NMT [32], to create an unsupervised EEG pre-training dataset that we refer to as EEGComb2. The resulting EEGComb2 dataset for self-supervised pertaining contains EEG recordings from 4438 subjects. We pretrain EEG-VJEPA model with the vision transformer (ViT) as the encoder´s backbone on EEGComb2 dataset. We train a ViT-B/$h\times w\times t$, a ViT-M/$h\times w\times t$, and a ViT-S/$h\times w\times t$ transformer models with various tublet sizes. We use different batch sizes (4,5,6,10,30,40) during experiments; see the combination of different hyper-parameters during pre-training in detail in Table 1. Key hyperparameters include:
@@
Baselines: Our self-supervised baseline consists of EEG2REP [19], LaBraM [20], and the contrastive learning (CL) based model [21], both trained on the TUAB dataset [34]. The EEG2REP model is based on the I-JEPA [30] model, while the contrastive learning model is based on the concept of multi-view learning [35] with ViT as its backbone. The LaBraM [20] is also a transformer based model pre-trained over 20 EEG datasets by vector quantized neural spectrum prediction, where the tokenizer predicts the Fourier spectrum of the original signal. Additionally, we benchmark against supervised methods, including BSVT [36], a vision transformer-based approach, and Chrononet [37], an RNN-based model.
@@
Figure 5: EEG-VJEPA locates regions of interest using spatiotemporal tokens. We roll out attention weights through all layers to visualize the attention along the EEG channel and time dimensions in 2D. The corresponding PSD plots show how EEG signal power is distributed across frequency bands, highlighting differences between normal and abnormal samples, across three different subjects in each class.
@@ 4.2 Evaluation on TUAB dataset
Fine-Tuning: We evaluate the EEG-VJEPA model by fine-tuning on a subset of the TUAB dataset as mentioned in Section 4.2. The smaller version of the EEG-VJEPA model with ViT-M/$4\times 30\times 4$ achieves an accuracy of 85.80%±$\pm{0.5}$, F1 score of 85.6%±$\pm{0.4}$, and an AUROC of 88.5%±$\pm{0.2}$ under fine-tuning settings. It outperforms the state-of-the-art baseline models and is on par with Chrononet, a model trained under a fully supervised setting. We also include in the comparison the EEG-VJEPA model with ViT-B/$4\times 30\times 2$ backbone (see Table 3 and Table 4 ).
EEG-VJEPA learns meaningful representations: To observe different structures learned by EEG-VJEPA, we uncover the learned embeddings and map their relationship with pathological and demographic attributes such as gender (male/female), age groups. The 384-dimensional embeddings obtained on the TUAB dataset are projected onto a two-dimensional space with Uniform Manifold Approximation [39]. In Figure 4(a), we observe age associated local structures with a gradient. The younger age groups tend to cluster towards left region of the plot and the older age groups lean more towards the right, although with some overlap. Similarly, Figure 4(b) reveals a global structure based on pathology. While observing Figure 4(a) and Figure 4(b) simultaneously, we can observe an alignment where normal EEG samples more often corresponded with younger age groups region and abnormal samples with older groups region in the plot. The gender-related global structure in Figure 4(c) shows a distribution with some overlap between male and female groups. This suggests that the model captured gender-specific features, however, there are shared physiological characteristics in the EEG signals across genders. It is worth noting that factors such as dimensionality reduction, generalized features, and the presence of noise may also contribute to the variability and overlap in the observed clusters.
@@
As shown in Figure 5, the 2D attention maps demonstrate the model’s ability to focus on spatiotemporal regions of interest within the input sequences. By propagating the attention through all layers, we visualize the dynamic focus of the model on diagnostically relevant spatial channels and temporal frames. The power spectral density (PSD) plots show how EEG signal power is distributed across different frequency bands, highlighting spectral differences between normal and abnormal samples that can help in distinguishing brain states. We computed each band’s relative power from the PSD—guided by the model’s attention maps—for both normal and abnormal recordings. In the normal condition, beta rhythms accounted for 27.81% of total power; in the abnormal condition, beta power fell to 11.48%. This decrease in faster oscillations echoes prior reports of diminished beta activity in pathological EEGs [42].
@@ 5 Discussion
EEG signals exhibit unique spatial and temporal characteristics that provide critical insights into dynamic brain activity, essential for effective EEG analysis and BCI applications. To address these challenges, we proposed EEG-VJEPA, a self-supervised model designed to learn robust representations of EEG data. Pre-trained on two diverse datasets, the proposed model consistently outperforms baseline methods, notably surpassing the state-of-the-art LaBraM [20], which was trained on 20 different EEG datasets, and achieving performance on par with the fully supervised ChronoNet model [37]. It also exhibits a substantial performance advantage over self-supervised approaches such as contrastive learning [21] and the JEPA-based EEG2REP model [19], underscoring the benefit of video-inspired joint embedding predictive modeling for EEG signals.
@@ 6 Conclusion
In this work, we introduced EEG-VJEPA, a novel self-supervised learning framework for EEG representation learning, which adapts the Video Joint Embedding Predictive Architecture (V-JEPA) to model multichannel EEG as spatiotemporal sequences. By leveraging masked latent prediction and a Vision Transformer backbone, EEG-VJEPA learns semantically meaningful and structured representations without the need for expert-labeled data. We pre-trained the model on two large-scale public EEG datasets and evaluated it on the Temple University Hospital Abnormal EEG Corpus (TUAB), a clinically realistic and heterogeneous benchmark. EEG-VJEPA outperforms prior self-supervised approaches and performs comparably to fully supervised models such as Chronet. Importantly, the model’s attention patterns and latent embeddings align with physiologically and clinically meaningful EEG features, offering a pathway to interpretability and trust.
@@
Finally, expanding EEG-VJEPA toward multimodal learning—combining EEG with speech, imaging, or other physiological signals—could open new avenues for holistic neurocognitive modeling and cross-modal representation learning.
Prospect of application. EEG-VJEPA provides a scalable and interpretable foundation model for EEG analysis, capable of learning clinically relevant features from unlabeled data. Its strong performance on a noisy, real-world dataset and alignment with neurophysiological patterns suggest applicability in clinical workflows such as EEG triage, Intensive Care Unit (ICU) monitoring, and early disease screening, such as dementia detection. With further validation and integration into clinician-facing systems, EEG-VJEPA could support human-AI collaboration in neurological diagnostics and improve access to high-quality EEG interpretation in under-resourced settings.

### v4

@@ Abstract
EEG signals capture brain activity with high temporal but low spatial resolution, supporting applications such as neurological diagnosis, cognitive monitoring, and brain-computer interfaces. However, effective analysis remains challenging due to limited labeled data, high dimensionality, and the lack of scalable models that fully capture spatiotemporal dependencies. Existing self-supervised learning (SSL) methods often focus on either spatial or temporal features in isolation, leading to suboptimal representations. To this end, we propose EEG-VJEPA, a novel adaptation of the Video Joint Embedding Predictive Architecture (V-JEPA) for EEG classification. By treating EEG as video-like sequences, EEG-VJEPA learns semantically meaningful spatiotemporal representations using joint embeddings and adaptive masking. To our knowledge, this is the first work that exploits V-JEPA for EEG classification and explores the visual concepts learned by the model. EEG-VJEPA achieves state-of-the-art performance on the publicly available Temple University Hospital (TUH) Abnormal EEG dataset, outperforming both self-supervised and fully supervised baselines. Likewise, we demonstrate the model’s good generalization ability on an independent, smaller clinical dataset from the General Hospital of Thessaloniki, involving dementia classification. Beyond classification, EEG-VJEPA captures physiologically meaningful spatiotemporal patterns, offering interpretable embeddings useful for human-AI collaboration in clinical workflows. These results highlight its potential as a scalable and trustworthy framework for real-world EEG analysis.
@@ 1 Introduction
    We present EEG-VJEPA, a novel self-supervised model based on the joint embedding predictive architecture that incorporates the spatial and temporal attributes of EEG signals. We treat EEG signals as video-like sequences, enabling the application of V-JEPA’s spatiotemporal masking strategy.
@@
    We demonstrate that self-supervised learning with the VJEPA strategy yields substantial performance gains over EEG2REP [19], LaBraM [20], and contrastive learning-based models [21] on the clinically relevant TUAB dataset, driven by the rich spatiotemporal representations learned in the latent space. Furthermore, we validate the generalization ability of EEG-VJEPA on an independent, smaller clinical dataset from the General Hospital of Thessaloniki.
@@
The remainder of the paper is structured as follows: Section 2 provides an overview of the SSL literature. In Section 3, we present the methodology of our work. Section 4 describes the experimental settings, highlighting the implementation details and results. In Section 5, we discuss our findings and implications. Finally, in section 6, we conclude the paper by discussing its limitations and outlining potential directions for future research.
@@ 2 Related Work
Substantial progress has been made in EEG representation learning, with advances in feature extraction, pre-training, predictive modeling, and downstream task performance. Early works transformed time-series EEG data into frequency domains and combined statistical and spectral features for classification using models such as SVMs [23]. Other studies leveraged joint representations from time and frequency domains via convolutional layers [24] or multi-view approaches such as wavelet transforms and attention-based gated recurrent units [25]. Incorporating brain connectivity measures like phase lag index and power spectral density has also enhanced predictive modeling [26].
Self-supervised learning (SSL) methods have shown promise in addressing the scarcity of labeled EEG data. Multi-view SSL approaches have fused diverse representations, such as time, frequency, and spatial features, for tasks like motor imagery classification and seizure prediction [27, 24]. A work by [21] developed a self-supervised approach for EEG representation learning using multi-view contrastive learning. The method utilized both time and frequency domain views of EEG data with data augmentation. [28] introduced EEGformer, a model combining convolutional and Transformer layers to capture temporal and regional EEG features. These methods demonstrated strong performance in emotion recognition and seizure detection. Recent works also highlight the potential of large-scale pre-training for EEG. Chen et al. [29] developed a 1.7TB EEG foundation model using a vector-quantized Transformer, offering interpretable representations with strong transfer learning capabilities. The Joint Embedding Predictive Architecture (JEPA), a self-supervised framework for learning semantic representations by predicting missing information in the latent space without data augmentation, was introduced by [30]. This approach has been adapted for various data types, including EEG signals. For instance, Guetschel et al. proposed Signal-JEPA which incorporates a domain-specific spatial block masking strategy to enhance cross-dataset transfer in EEG signal processing [31]. However, these approaches often fail to capture the long-range temporal context critical for EEG analysis. Inspired by V-JEPA [22], we extend this architecture to EEG by treating signals as video-like sequences.
@@ 3 Methodology
We process the EEG sequence by first dividing frames into non-overlapping spatiotemporal patches, also referred to as volumes or tubelets. Each patch captures a small region of the EEG signal across several channels and time points. We then apply a patch embedding layer, combining convolutional operations with linear projections, to map each patch to a fixed-dimensional embedding vector.
The X-encoder processes the sequence of patch embeddings after masking a portion of the patches. The masking strategy is crucial for creating the prediction task. We adapt V-JEPA’s multi-block masking approach, which involves masking large, spatially contiguous blocks that span the entire temporal dimension of the sequence. This strategy encourages the model to learn long-range spatiotemporal dependencies. During masking, the aspect ratio for the sampled blocks is randomly chosen in a certain range (see Table 1).
@@ 4.1 Experimental Setting
Datasets: We use three publicly available datasets, namely TUAB [18], NMT [32], and the EEG dataset from the General Hospital of Thessaloniki.
@@
General Hospital of Thessaloniki EEG dataset contains resting-state EEG recordings collected at the General Hospital of Thessaloniki [33]. Each recording is labeled according to clinical diagnosis, including Alzheimer’s disease (AD), frontotemporal dementia (FTD), and cognitively normal controls (CN). EEG data were acquired using a 19-channel cap, arranged according to the international 10–20 electrode placement system, with a referential montage referenced to Cz. All recordings were sampled at 500 Hz under resting-state, eyes-open conditions. The dataset comprises recordings from 36 individuals with AD, 23 with FTD, and 29 healthy controls
Preprocessing: We apply some basic preprocessing steps on the EEG signals to prepare them as input for our model. We crop the recordings to a fixed 5-minute time length and select a common subset of 19 channels. We downsampled the signals to 100Hz. We then apply bandpass filtering between 1-40 Hz. Afterward, we use a sliding window technique to produce 5-second overlapping windows of size $(19\times 500)$, transforming the EEG signal into a 3D tensor of size $118\times 19\times 500$. As the last step, we normalize the EEG data channel-wise to achieve a zero mean and unit standard deviation.
@@
Following the literature, we use the following naming scheme for the ViT models: ViT-Base (ViT-B, $L=12$, $NH=12$, $d=768$) with 85 M parameters, ViT-Medium (ViT-M, $L=12$, $NH=6$, $d=384$) with 21 M parameters, and ViT-Small (ViT-S, $L=12$, $NH=3$, $d=192$) with 5 M parameters, where $L$ is the number of transformer layers, each with a self-attention block of $NH$ heads. To indicate the tubelet size while patching, we extend the naming scheme, for example, ViT-B/$4\times 30\times 2$ denotes a ViT-Base backbone with the tubelet size of ($h\times w\times t=4\times 30\times 2$). We tested various ViT encoder sizes (base, medium, small) to evaluate the impact of model capacity on representation learning; refer to Table 1 for more details. In all our experiments, the tubelet height ($h$) and width($w$) differ due to the shape of our EEG input signals´ height and width ($19\times 500$). We also probe with the time dimension $t$ while patching.
Pre-training: We combine two publicly available datasets, namely TUH [18] and NMT [32], to create an unsupervised EEG pre-training dataset that we refer to as EEGComb2. The resulting EEGComb2 dataset for self-supervised pertaining contains EEG recordings from 4438 subjects. We pretrain EEG-VJEPA model with the vision transformer (ViT) as the encoder´s backbone on EEGComb2 dataset. We train a ViT-B/$h\times w\times t$, a ViT-M/$h\times w\times t$, and a ViT-S/$h\times w\times t$ transformer models with various tublet sizes. We use different batch sizes (4,5,6,10,30,40) during experiments; see the combination of different hyperparameters during pre-training in detail in Table 1. Key hyperparameters include:
@@
Baselines: Our self-supervised baseline consists of EEG2REP [19], LaBraM [20], and the contrastive learning (CL) based model [21], both trained on the TUAB dataset [35]. The EEG2REP model is based on the I-JEPA [30] model, while the contrastive learning model is based on the concept of multi-view learning [36] with ViT as its backbone. The LaBraM [20] is also a transformer based model pre-trained over 20 EEG datasets by vector quantized neural spectrum prediction, where the tokenizer predicts the Fourier spectrum of the original signal. Additionally, we benchmark against supervised methods, including BSVT [37], a vision transformer-based approach, and Chrononet [38], an RNN-based model. We also include a baseline SVM classifier [39], trained on a combination of time-domain features, frequency-domain features, connectivity metrics, and complexity-based measures, in our evaluation on the smaller, independent dataset from the General Hospital of Thessaloniki.
@@
Figure 5: EEG-VJEPA locates regions of interest using spatiotemporal tokens. We roll out attention weights through all layers to visualize the attention along the EEG channel and time dimensions in 2D. The corresponding PSD plots show how EEG signal power is distributed across frequency bands, highlighting differences between normal (left side) and abnormal (right side) samples, across five different subjects in each class.
@@ 4.2 Evaluation on TUAB dataset
Fine-Tuning: We evaluate the EEG-VJEPA model by fine-tuning on a subset of the TUAB dataset as mentioned in Section 4.2. The medium version of the EEG-VJEPA model with ViT-M/$4\times 30\times 4$ achieves an accuracy of 85.80%±$\pm{0.5}$, F1 score of 85.6%±$\pm{0.4}$, and an AUROC of 88.5%±$\pm{0.2}$ under fine-tuning settings. It outperforms the state-of-the-art baseline models and is on par with Chrononet, a model trained under a fully supervised setting. We also include in the comparison the EEG-VJEPA model with ViT-B/$4\times 30\times 2$ backbone (see Table 3 and Table 4 ).
EEG-VJEPA learns meaningful representations: To observe different structures learned by EEG-VJEPA, we uncover the learned embeddings and map their relationship with pathological and demographic attributes such as gender (male/female), age groups. The 384-dimensional embeddings obtained on the TUAB dataset are projected onto a two-dimensional space with Uniform Manifold Approximation [41]. In Figure 4(a), we observe age-associated local structures with a gradient. The younger age groups tend to cluster towards the left region of the plot, and the older age groups lean more towards the right, although with some overlap. Similarly, Figure 4(b) reveals a global structure based on pathology. While observing Figure 4(a) and Figure 4(b) simultaneously, we can observe an alignment where normal EEG samples more often corresponded with younger age groups region and abnormal samples with the older groups region in the plot. The gender-related global structure in Figure 4(c) shows a distribution with some overlap between male and female groups. This suggests that the model captured gender-specific features, however, there are shared physiological characteristics in the EEG signals across genders. It is worth noting that factors such as dimensionality reduction, generalized features, and the presence of noise may also contribute to the variability and overlap in the observed clusters.
@@
As shown in Figure 5, the 2D attention maps demonstrate the model’s ability to focus on spatiotemporal regions of interest within the input sequences. By propagating the attention through all layers, we visualize the dynamic focus of the model on diagnostically relevant spatial channels and temporal frames. The power spectral density (PSD) plots show how EEG signal power is distributed across different frequency bands, highlighting spectral differences between normal and abnormal samples that can help in distinguishing brain states. We computed each band’s relative power from the PSD, guided by the model’s attention maps, for both normal and abnormal recordings. In the normal condition, beta rhythms accounted for 27.81% of total power; in the abnormal condition, beta power fell to 11.48%. This decrease in faster oscillations echoes prior reports of diminished beta activity in pathological EEGs [44].
@@
4.3 Evaluation on the General Hospital of Thessaloniki EEG Dataset
To assess the generalization capabilities of EEG-VJEPA on a smaller, independent dataset, we utilized the General Hospital of Thessaloniki EEG dataset (88 subjects), comprising individuals diagnosed with Alzheimer’s disease (AD), frontotemporal dementia (FTD), and cognitively normal controls (CN). For binary classification, we merged the AD and FTD groups into a single dementia category.
Following the evaluation method shown in Figure 3, we fine-tune EEG-VJEPA on the General Hospital of Thessaloniki EEG dataset. The EEG-VJEPA model based on ViT-M/$4\times 30\times 4$ achieves an accuracy of 83.34%±$\pm{0.2}$, F1 score of 83.48%±$\pm{0.2}$ under fine-tuning settings. While the performance is strong, it remains slightly below the baseline SVM model (see Table 5). The performance gap likely reflects the small dataset size and the handcrafted features tailored to dementia in the SVM approach. In contrast, our method requires no manual feature design and learns general EEG patterns through pretraining, offering a flexible foundation that may improve with larger or more diverse pretraining data.
Table 5: Classification performance of EEG-VJEPA on the General Hospital of Thessaloniki EEG dataset (88 subjects). Results are reported for the ViT-M/$4\times 30\times 4$ configuration under fine-tuning and frozen settings.
Method Accuracy Recall F1
SVM [39] 93.5% 90.0% -
ViT-M/$4\times 30\times 4$ 67.00%±$\pm{0.2}$ 54.66%±$\pm{0.2}$ 53.21%±$\pm{0.2}$
ViT-M/$4\times 30\times 4$ 83.34%±$\pm{0.2}$ 83.70%±$\pm{0.2}$ 83.48%±$\pm{0.2}$
Frozen evaluation, Fine-tuning.
@@ 5 Discussion
EEG signals exhibit unique spatial and temporal characteristics that provide critical insights into dynamic brain activity, essential for effective EEG analysis and BCI applications. To address these challenges, we proposed EEG-VJEPA, a self-supervised model designed to learn robust representations of EEG data. Pre-trained on two diverse datasets, the proposed model consistently outperforms baseline methods, notably surpassing the state-of-the-art LaBraM [20], which was trained on 20 different EEG datasets, and achieving performance on par with the fully supervised ChronoNet model [38]. It also exhibits a substantial performance advantage over self-supervised approaches such as contrastive learning [21] and the JEPA-based EEG2REP model [19], underscoring the benefit of video-inspired joint embedding predictive modeling for EEG signals. We further evaluated EEG-VJEPA on an independent, smaller clinical dataset from the General Hospital of Thessaloniki, focusing on the binary classification of dementia versus healthy controls. Despite the limited sample size, EEG-VJEPA demonstrated strong generalization capabilities and competitive performance compared to the supervised baseline, highlighting its potential applicability in real-world clinical scenarios where labeled data are scarce.
@@ 6 Conclusion
In this work, we introduced EEG-VJEPA, a novel self-supervised learning framework for EEG representation learning, which adapts the Video Joint Embedding Predictive Architecture (V-JEPA) to model multichannel EEG as spatiotemporal sequences. By leveraging masked latent prediction and a Vision Transformer backbone, EEG-VJEPA learns semantically meaningful and structured representations without the need for expert-labeled data. We pre-trained the model on two large-scale public EEG datasets and evaluated it on the Temple University Hospital Abnormal EEG Corpus (TUAB), a clinically realistic and heterogeneous benchmark. EEG-VJEPA outperforms prior self-supervised approaches and performs comparably to fully supervised models such as Chronet. Moreover, it demonstrates strong generalization to a smaller, independent dataset from the General Hospital of Thessaloniki. Importantly, the model’s attention patterns and latent embeddings align with physiologically and clinically meaningful EEG features, offering a pathway to interpretability and trust.
@@
Finally, expanding EEG-VJEPA toward multimodal learning, combining EEG with speech, imaging, or other physiological signals, could open new avenues for holistic neurocognitive modeling and cross-modal representation learning.
Prospect of application. EEG-VJEPA provides a scalable and interpretable path towards a foundation model for EEG analysis, capable of learning clinically relevant features from unlabeled data. Its strong performance on a noisy, real-world dataset and alignment with neurophysiological patterns suggest applicability in clinical workflows such as EEG triage, Intensive Care Unit (ICU) monitoring, and early disease screening, such as dementia detection. With further validation and integration into clinician-facing systems, EEG-VJEPA could support human-AI collaboration in neurological diagnostics and improve access to high-quality EEG interpretation in under-resourced settings.